A four- lane road need to mix a liquid asphalt with a soil is (24 cm) thick and liquid
asphalt should be seven percent by weight of soil, if the required quantity of the
material per one Km is (569225) Liters with S.G. is (0.85) at spraying temperature.
Find max. dry density of soil.
Answer #1

Given,

Four lane road, assume width = 14 m

Quantity of asphalt per 1 km = 569225 litres = 569.225 3

S.G = 0.85

Density of asphalt = 0.85 * 1000 = 850 Kg/m^{3}

Weight of asphalt per 1 km = density * volume = 850 * 569.225 = 483,841.25 kg

Weight of asphalt is 7% Weight of soil

Weight of soil = Weight of asphalt/0.07 = 483,841.25/0.07 = 6,912,018 kg

Volume of soil per 1 km = 1000 * 14 * 0.24 = 3,360 3

Dry density of soil = Weight/Volume = 6,912,018/3,360 = 2057.15 Kg/m^{3}